If you add 4 to my age and divide the sum by 4, you will have my son's age. But 6 years ago I was 7.5 times as old as my son. Find our ages
Answer by lwsshak3(11628)

If you add 4 to my age and divide the sum by 4, you will have my son's age. But 6 years ago I was 7.5 times as old as my son. Find our ages.
***
let x= father's age
(x+4)/4=son's age
..
6 yrs ago:
x-6= father's age
6yrs ago:
x-6= father's age
(x+4)/4)-6=son's age
x-6=7.5((x+4)/4)-6)
x-6=(7.5x+30)/4-45
lcd:4
4x-24=7.5x+30-180
3.5x=126
x=36
(x+4)/4=40/4=10
father's age=36
son's age=10